solve the equation
6t + 5(t-1)/3= 5
exact answer
t=?

Answer by TimothyLamb(4379)

6t + 5(t - 1)/3 = 5
6t + (5/3)t - 5/3 = 5
(18/3)t + (5/3)t = 15/3 + 5/3
(23/3)t = 20/3
t = (20/3)*(3/23)
t = (20/23)
t = 0.869565217391
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---